# A fast particle-field method for 3D Chemotaxis systemsJingyuan Hu, Zhongjian Wang, Jack Xin and Zhiwen Zhang

### v2

@@ 4.1 Preliminary Estimates and Main Convergence Proof
for some constant $M_{2}$, and can achieve the one-step error control
@@
for some constant $K$. Then, the final mean absolute error $d^{\left(N_{T}\right)}$ satisfies
@@
From the growth estimate hypothesis,
@@
By the induction hypothesis and one-step error control $F^{(k)}\leq G^{(k)}+\\ K\max(d^{(0)},\ldots,d^{(k)})$, we derive
@@
The mean absolute error $d^{(n)}$ is controlled by $d^{(n-1)}$ with probability at least $1-e^{-\frac{S_{1}^{2}}{2}}$:
@@
where $M_{3},N_{2}$ are defined in Assumption 4.5, and $S_{1}>0$ is a parameter controlling the probability bound, $\nabla c^{(n-1)}=\nabla c(\cdot,(n-1)\tau)$.
@@
Let the drift mapping $g^{(n-1)}(\mathbf{X})=\mathbf{X}+\chi\tau\nabla c^{(n-1)}(\mathbf{X})$, $\overline{g}^{(n-1)}(\mathbf{X})=\mathbf{X}+\chi\tau\nabla\hat{c}^{(n-1)}(\mathbf{X})$ and $\hat{g}^{(n-1)}(\mathbf{X})=\mathbf{X}+\chi\tau I^{*}\nabla\hat{c}^{(n-1)}(\mathbf{X})$.
Since the updates of the tensors $\mathbf{X}^{(n)}$ and $\widehat{\mathbf{X}}^{(n)}$ use the same diffusion $\mathbf{W}^{(n)}$, we have by triangle inequality,
@@
For the estimate of the first part, notice that $\widetilde{\mathbf{X}}_{p}^{(n-1)}\sim\text{i.i.d.}\ \mathbf{X}^{(n-1)}$. The expectation is given by
@@
and the variance has an upper bound
@@
$\leq\frac{\chi^{2}\tau^{2}}{M_{0}}\int_{\Omega}4M_{1}^{2}\rho^{(n-1)}\mathrm{d}\mathbf{x}$
$=\chi^{2}\tau^{2}\cdot 4M_{1}^{2}.$
@@
with probability over $1-e^{-\frac{S_{1}^{2}}{2}}$.
For the second part, we employ Assumption 4.5 that the second derivatives of $\hat{c}$ are bounded by $M_{2}$. Note the displacements $\epsilon_{p}^{(n-1)}:=\widehat{\mathbf{X}}_{p}^{(n-1)}-\widetilde{\mathbf{X}}_{p}^{(n-1)}$. Then $d^{(n)}=\frac{1}{P}\sum\limits_{p=1}^{P}|\epsilon_{p}^{(n)}|$, and
@@
By Lemma 3.11, the last term of the decomposition admits the bound:
@@
Combining the three estimates through (70), (71), and (72), we obtain the target estimate.
@@
The 2-norm of the difference between the auxiliary gradient $\nabla\tilde{c}$ and the empirical gradient $\nabla\hat{c}$ is limited by
@@
where $D^{(n)}=\max(d^{(0)},\ldots,d^{(n-1)})$ denotes the maximum historical error across iterations, $T$ is the total time, and the interpolation error constant $C=\frac{513}{16}$ for fourth-order Particle-to-Grid.
@@
We invoke the Parseval theorem to establish the required bound:
@@
where the truncated frequency index set $U_{0}:=\left\{-\frac{H_{0}}{2},\ldots,\frac{H_{0}}{2}\right\}^{3}.$
The dynamics of spectral error $e^{(n)}$ derive directly from the update rules (10) and (48). For any frequency index $\mathbf{q}\in U_{0}$, the error evolution satisfies:
@@
where $\tilde{\beta}_{\mathbf{q}}:=\mathcal{F}_{H}\tilde{\rho}(\mathbf{q}\frac{2\pi}{L}),$ $\hat{\beta}_{\mathbf{q}}:=\mathcal{F}_{H}\hat{\rho}(\mathbf{q}\frac{2\pi}{L})$ and $\overline{\beta}_{\mathbf{q}}:=\mathcal{F}_{H}I\left(\hat{\rho}\right)(\mathbf{q}\frac{2\pi}{L})$ denote the discrete Fourier coefficients of the density fields.
Due to Lemma 3.9, we have the first part
@@
Recall the displacements $\epsilon_{p}^{(n-1)}:=\widehat{\mathbf{X}}_{p}^{(n-1)}-\widetilde{\mathbf{X}}_{p}^{(n-1)}$, we have the second part
(77) $\bigl|\tilde{\beta}_{\mathbf{q}}^{(n-1)}-\hat{\beta}_{\mathbf{q}}^{(n-1)}\bigr|\leq\frac{M_{0}}{PL^{3}}\left|e^{-i\widetilde{\mathbf{X}}_{1}^{(n-1)}\mathbf{q}\frac{2\pi}{L}}\epsilon_{1}+\cdots+e^{-i\widetilde{\mathbf{X}}_{P}^{(n-1)}\mathbf{q}\frac{2\pi}{L}}\epsilon_{P}\right|\left|\mathbf{q}\frac{2\pi}{L}\right|,$
Note the spectral displacement error $\nu_{\mathbf{q}}^{(n)}:=\frac{1}{PL^{3}}\sum_{p=1}^{P}e^{-i\widetilde{\mathbf{X}}_{p}^{(n)}\mathbf{q}\frac{2\pi}{L}}\epsilon_{p}^{(n)}$. The error bound arises from interpolation errors (noted $\hat{e}_{\mathbf{q}}^{(n)}$) and displacement-induced errors (noted $\check{e}_{\mathbf{q}}^{(n)}$):
(78) $|e_{\mathbf{q}}^{(n)}|\leq|\hat{e}_{\mathbf{q}}^{(n)}|+|\check{e}_{\mathbf{q}}^{(n)}|,$
@@
For (79), summing the decaying geometric series yields:
@@
For (80), note the single-step decay factor $\kappa_{\mathbf{q}}=\frac{1}{1+\frac{\tau}{\epsilon}\left(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}\right)}$, we have
@@
Applying the spectral error bounds through the Parseval equivalence (74), we translate the frequency-domain estimates back to physical space to conclude the proof.
@@
By the Parseval theorem, the single-step displacement square norm $\sum_{\mathbf{q}\in U_{0}}\left|\nu_{\mathbf{q}}^{(s)}\right|^{2}=O\left(L^{-3}(d^{(s)})^{2}\right)$, and since the step number $n\leq N_{T}=\frac{T}{\tau}$, the accumulated displacement square norm $\sum_{0\leq s<n}\sum_{\mathbf{q}\in U_{0}}\left|\nu_{\mathbf{q}}^{(s)}\right|^{2}=O\left(N_{T}L^{-3}(D^{(n)})^{2}\right)$. Substituting (83)-(85) into (74) yields the desired result.
@@
The spectral displacement error $\nu_{\mathbf{q}}^{(n)}=\frac{1}{PL^{3}}\sum_{p=1}^{P}e^{-i\widetilde{\mathbf{X}}_{p}^{(n)}\mathbf{q}\frac{2\pi}{L}}\epsilon_{p}^{(n)},$ is a tensor of size $H\times H\times H\times 3$, and the Fourier transform is applied on the first three dimensions.
@@
The 2-norm of the difference between the reference gradient $\nabla c$ and the auxiliary gradient $\nabla\tilde{c}$ is limited by
@@
where $N_{1,2}$ is defined in Assumption 4.5 and the random space is mentioned in Remark 4.2.
@@
By the Parseval theorem, notice that $\tilde{\alpha}^{(n)}_{\mathbf{q}}=0$ for $q\in\mathbb{Z}^{3}\setminus U_{0}$, we have
@@
For high frequencies $\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}$, the error $e^{(n)}_{\mathbf{q}}=\alpha^{(n)}_{\mathbf{q}}$. The update of $\alpha_{\mathbf{q}}$ satisfies
@@
By the inequality $(u+v)^{2}\leq(1+s)u^{2}+\left(1+\frac{1}{s}\right)v^{2}$ setting $s=\frac{\tau}{\epsilon}\left(\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}+k^{2}\right),\\ \quad u=\frac{\left|\alpha^{(n-1)}_{\mathbf{q}}\right|}{1+s},\quad v=\frac{\left|\beta^{(n-1)}_{\mathbf{q}}\right|}{\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}+k^{2}+\frac{\epsilon}{\tau}}$, we get the following bound:
@@
For all $\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}$, multiplying by $|\mathbf{q}\frac{2\pi}{L}|^{2}$ and summing yields:
@@
Summing over $\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}$:
(91) $\sum_{\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}}\left|\mathbf{q}\frac{2\pi}{L}\alpha^{(n)}_{\mathbf{q}}\right|^{2}$
$\leq\sum_{\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}}\biggl(\frac{1}{1+\frac{\tau}{\epsilon}\left((H_{0}\frac{2\pi}{L})^{2}+k^{2}\right)}\left|\mathbf{q}\frac{2\pi}{L}\alpha^{(n-1)}_{\mathbf{q}}\right|^{2}+\frac{\left|\beta^{(n-1)}_{\mathbf{q}}\right|^{2}}{|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}+\frac{\epsilon}{\tau}}\biggr).$
Similar to (81), this reduces to:
@@
For the low-frequency terms, the independence of particles $\widetilde{\mathbf{X}}_{1}^{(n)},\ldots,\widetilde{\mathbf{X}}_{P}^{(n)}\sim\mathbf{X}^{(n)}$ implies:
@@
Applying the attenuation factor $\frac{1}{|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}}$, we get
@@
The factor $6H_{0}$ originates from a spherical shell summation. For each $1\leq h\leq H_{0}$, there exist $6h^{2}$ frequency indices $\mathbf{q}$ with the $L^{1}$ norm $|\mathbf{q}|_{1}=h$. The $h^{-2}$ scaling of the spectral terms $\frac{|\mathbf{q}\frac{2\pi}{L}|^{2}}{(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2})^{2}}$ cancels this growth, leaving a linear accumulation $H_{0}$.
Combining high-frequency errors (92) and low-frequency variance (94) through the Parseval decomposition (87), we establish the final error bound.
The error of the Euler–Maruyama semi‑discretization over the time interval is stated as follows (see [milstein2004stochastic, p. 15]).
@@
Let $\{\mathbf{X}_{p}\}_{1\leq p\leq P}$ be $P$ independent realizations of the stochastic differential equation (11), and the Brownian increments $\mathbf{W}_{p}$ satisfy
@@
The Euler-Maruyama semi-discretization has a strong order of accuracy $\eta=1$. More precisely, for $t=n\tau\leq T$ with integer $n$,
@@
where $C_{0}=C_{0}(J_{1},M_{1},M_{2},M_{3})$ is independent of $\tau$ and $t$.
The main convergence results are presented as follows.
@@
For SIPF-PIC using fourth-order Particle-to-Grid and second-order Grid-to-Particle interpolation, the growth rate of the mean absolute error $\hat{d}^{(n)}:=\frac{1}{P}\sum_{1\leq p\leq P}|\mathbf{X}_{p}(n\tau)-\widehat{\mathbf{X}}_{p}^{(n)}|$ at final time $T$ satisfies
@@
$+\left.\left.C_{3}L^{2}H_{0}^{-2}+M_{3}L^{2}H^{-2}+2S_{1}M_{1}P^{-\frac{1}{2}}\right)\right)$
with probability $1-N_{T}e^{-\frac{S_{1}^{2}}{2}}$, where $\chi$ is chemotactic sensitivity in system (1), $K$, $C_{0}$, $C_{1}$, $C_{2}$, $C_{3}$ and $S_{1}$ are proof-derived constants independent of the period $L$, the grid resolution $H$, the cutoff threshold $H_{0}$, the number of particles $P$ and the timestep $\tau$.
@@
First, we employ the Discrete Gronwall Inequality (Lemma 4.6) to derive an estimate for $d^{(n)}$. From Lemma 4.8, we get:
@@
We decompose the gradient error terms using the auxiliary-empirical and reference-auxiliary bounds from Lemmas 4.10 and 4.13. First, the gradient discrepancy is split via the inequality:
@@
Substituting the reference-auxiliary expectation bound (86) and the auxiliary-
empirical bound (73) into the Gronwall framework yields the composite error growth factor $K=\sqrt{2TN_{2}}$ and residual terms
@@
with $C_{1}=513\pi^{3}\sqrt{N_{2}},C_{2}=\sqrt{N_{2}},C_{3}=(2\pi)^{-2}\frac{\sqrt{N_{2}N_{1,2}}}{M_{0}}$ defined via the problem constants. The final error bound follows from the Gronwall factor $(1+(K+\chi M_{2})\tau)^{N_{T}}\leq e^{T(K+\chi M_{2})}$ and the accumulation of residuals, establishing the growth rate
@@
which follows from the triangle inequality and the Cauchy–Schwarz inequality, together with Lemma 4.15, we obtain the claimed convergence rate (97).
@@
By optimizing the cutoff threshold at $H_{0}=H^{\frac{8}{13}}$, the error bound simplifies to:
@@
with probability $1-N_{T}e^{-\frac{S_{1}^{2}}{2}}$.

### v3

@@ 4.1 Main convergence results
for some growth constant $M_{2}$ and one-step error $F^{(n)}$, and that we can achieve the one-step error control
@@
for some local truncation error $G^{(n)}$ and error amplification constant $K$. Then, the final mean absolute error $d^{\left(N_{T}\right)}$ satisfies
@@
From the growth estimate in the hypothesis,
@@
By the induction hypothesis and the one-step error control $F^{(k)}\leq G^{(k)}+K\max(d^{(0)},\\ \ldots,d^{(k)})$, we first substitute these into the inequality for $d^{(k+1)}$ and then simplify to derive
@@
The mean absolute error $d^{(n)}$ is controlled by $d^{(n-1)}$ with a probability of at least $1-e^{-\frac{S_{1}^{2}}{2}}$:
@@
where $M_{3}$ and $N_{2}$ are defined in Assumption 4.5, $S_{1}>0$ is a parameter controlling the probability bound, and $\nabla c^{(n-1)}=\nabla c(\cdot,(n-1)\tau)$.
@@
Define the mappings $g^{(n-1)}(\mathbf{X})=\mathbf{X}+\chi\tau\nabla c^{(n-1)}(\mathbf{X})$, $\overline{g}^{(n-1)}(\mathbf{X})=\mathbf{X}+\chi\tau\nabla\hat{c}^{(n-1)}(\mathbf{X})$, and $\hat{g}^{(n-1)}(\mathbf{X})=\mathbf{X}+\chi\tau I^{*}\nabla\hat{c}^{(n-1)}(\mathbf{X})$, respectively. Since the updates of the solutions $\mathbf{X}^{(n)}$ and $\widehat{\mathbf{X}}^{(n)}$ use the same diffusion $\mathbf{W}^{(n)}$, we have by the triangle inequality,
@@
For the estimate of the first part on the right-hand side of (59), note that $\widetilde{\mathbf{X}}_{p}^{(n-1)}$ are i.i.d. and follow the same distribution as $\mathbf{X}^{(n-1)}$. Its expectation is given by
@@
and its variance has an upper bound
@@
$\leq\frac{\chi^{2}\tau^{2}}{M_{0}}\int_{\Omega}4M_{1}^{2}\rho^{(n-1)}\mathrm{d}\mathbf{x}=\chi^{2}\tau^{2}\cdot 4M_{1}^{2}.$
@@
with probability at least $1-e^{-\frac{S_{1}^{2}}{2}}$.
For the second part on the right-hand side of (59), we employ Assumption 4.5, which states that the second derivatives of $\hat{c}$ are bounded by $M_{2}$. Define the displacements as $\epsilon_{p}^{(n-1)}:=\widehat{\mathbf{X}}_{p}^{(n-1)}-\widetilde{\mathbf{X}}_{p}^{(n-1)}$. Recall that $d^{(n)}=\frac{1}{P}\sum\limits_{p=1}^{P}|\epsilon_{p}^{(n)}|$, and
@@
By Lemma 3.11, the last term on the right-hand side of (59) admits the following bound
@@
Combining the three estimates via (62), (63), and (64), we obtain the target estimate in Lemma 4.8.
@@
The $2$-norm of the difference between the auxiliary gradient $\nabla\tilde{c}$ and the empirical gradient $\nabla\hat{c}$ over the domain $\Omega$ is bounded by
@@
where $D^{(n)}=\max(d^{(0)},\ldots,d^{(n-1)})$ denotes the maximum historical error across iterations, $T$ is the total computational time, and the interpolation error constant is $C=\frac{513}{16}$ for the fourth-order particle-to-grid interpolation.
@@
We use Parseval’s theorem to establish the required bound:
@@
where the truncated frequency index set is defined as $U_{0}:=\left\{-\frac{H_{0}}{2},\ldots,\frac{H_{0}}{2}\right\}^{3}.$
The dynamics of spectral error $e_{\mathbf{q}}^{(n)}$ defined in Definition 4.3 follow directly from the update formulas (10) and (41). For any frequency index $\mathbf{q}\in U_{0}$,
@@
where $\tilde{\beta}_{\mathbf{q}}:=\mathcal{F}_{H}\tilde{\rho}(\mathbf{q}\frac{2\pi}{L})$, $\hat{\beta}_{\mathbf{q}}:=\mathcal{F}_{H}\hat{\rho}(\mathbf{q}\frac{2\pi}{L})$, and $\overline{\beta}_{\mathbf{q}}:=\mathcal{F}_{H}I\left(\hat{\rho}\right)(\mathbf{q}\frac{2\pi}{L})$ denote the discrete Fourier coefficients of the respective density fields.
By Lemma 3.9, we have the following bound for the first term:
@@
Recall the displacement definition $\epsilon_{p}^{(n-1)}:=\widehat{\mathbf{X}}_{p}^{(n-1)}-\widetilde{\mathbf{X}}_{p}^{(n-1)}$. We have the following bound for the second term:
(69) $\bigl|\tilde{\beta}_{\mathbf{q}}^{(n-1)}-\hat{\beta}_{\mathbf{q}}^{(n-1)}\bigr|\leq\frac{M_{0}}{PL^{3}}\left|\sum_{p=1}^{P}e^{-i\widetilde{\mathbf{X}}_{p}^{(n-1)}\mathbf{q}\frac{2\pi}{L}}\epsilon_{p}^{(n-1)}\right|\left|\mathbf{q}\frac{2\pi}{L}\right|.$
We next define the spectral displacement error $\nu_{\mathbf{q}}^{(n)}:=\frac{1}{PL^{3}}\sum_{p=1}^{P}e^{-i\widetilde{\mathbf{X}}_{p}^{(n)}\mathbf{q}\frac{2\pi}{L}}\epsilon_{p}^{(n)}$, and the interpolation errors $\hat{e}_{\mathbf{q}}^{(n)}$ and displacement-induced errors $\check{e}_{\mathbf{q}}^{(n)}$ as
(70) $\hat{e}_{\mathbf{q}}^{(0)}=0,\hat{e}_{\mathbf{q}}^{(n)}=\frac{1}{1+\frac{\tau}{\epsilon}\left(\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}+k^{2}\right)}e^{(n-1)}_{\mathbf{q}}+\frac{1}{\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}+k^{2}+\frac{\epsilon}{\tau}}(\hat{\beta}^{(n-1)}_{\mathbf{q}}-\overline{\beta}^{(n-1)}_{\mathbf{q}})$,
$\check{e}_{\mathbf{q}}^{(0)}=0,\check{e}_{\mathbf{q}}^{(n)}=\frac{1}{1+\frac{\tau}{\epsilon}\left(\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}+k^{2}\right)}e^{(n-1)}_{\mathbf{q}}+\frac{1}{\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}+k^{2}+\frac{\epsilon}{\tau}}(\tilde{\beta}^{(n-1)}_{\mathbf{q}}-\hat{\beta}^{(n-1)}_{\mathbf{q}})$.
By (67) we have $|e_{\mathbf{q}}^{(n)}|\leq|\hat{e}_{\mathbf{q}}^{(n)}|+|\check{e}_{\mathbf{q}}^{(n)}|.$
These two error terms admit the following estimates, respectively
@@
For (71), summing the decaying geometric series yields
@@
For (72), we first define the single-step decay factor as $\kappa_{\mathbf{q}}=\frac{1}{1+\frac{\tau}{\epsilon}\left(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}\right)}$, and then derive the estimate as follows:
@@
Applying the above spectral error bounds and using Parseval’s identity (66), we translate the frequency-domain estimates back to the physical space to proceed with the proof. Specifically, we can obtain that
@@
By Parseval’s theorem, the single-step displacement square norm $\sum_{\mathbf{q}\in U_{0}}\left|\nu_{\mathbf{q}}^{(s)}\right|^{2}=O\left(L^{-3}(d^{(s)})^{2}\right)$. Since the number of time steps satisfies $n\leq N_{T}=\frac{T}{\tau}$, the accumulated displacement square norm is $\sum_{0\leq s<n}\sum_{\mathbf{q}\in U_{0}}\left|\nu_{\mathbf{q}}^{(s)}\right|^{2}=O\left(N_{T}L^{-3}(D^{(n)})^{2}\right)$. Finally, substituting (75)-(77) into (66) yields the desired result.
@@
The spectral displacement error $\nu_{\mathbf{q}}^{(n)}=\frac{1}{PL^{3}}\sum_{p=1}^{P}e^{-i\widetilde{\mathbf{X}}_{p}^{(n)}\mathbf{q}\frac{2\pi}{L}}\epsilon_{p}^{(n)},$ is a tensor of size $H\times H\times H\times 3$, with the Fourier transform applied to its first three dimensions.
@@
The $2$-norm of the difference between the reference gradient $\nabla c$ and the auxiliary gradient $\nabla\tilde{c}$ is bounded by
@@
where $N_{1,2}$ is defined in Assumption 4.5, and the underlying probability space is specified in Remark 4.2.
@@
By Parseval’s theorem and noting that $\tilde{\alpha}^{(n)}_{\mathbf{q}}=0$ for $q\in\mathbb{Z}^{3}\setminus U_{0}$, we have
@@
For high frequencies $\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}$, the error satisfies $e^{(n)}_{\mathbf{q}}=\alpha^{(n)}_{\mathbf{q}}$. The update formula for $\alpha_{\mathbf{q}}$ is given by
@@
Applying the inequality $(u+v)^{2}\leq(1+s)u^{2}+\left(1+\frac{1}{s}\right)v^{2}$ with $s=\frac{\tau}{\epsilon}\left(\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}+k^{2}\right)$, $u=\frac{\left|\alpha^{(n-1)}_{\mathbf{q}}\right|}{1+s}$, and $v=\frac{\left|\beta^{(n-1)}_{\mathbf{q}}\right|}{\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}+k^{2}+\frac{\epsilon}{\tau}}$, we obtain the following bound:
@@
For all $\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}$, multiplying through by $|\mathbf{q}\frac{2\pi}{L}|^{2}$ yields
@@
Summing over $\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}$ gives
(83) $\sum_{\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}}\left|\mathbf{q}\frac{2\pi}{L}\alpha^{(n)}_{\mathbf{q}}\right|^{2}\leq\sum_{\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}}\biggl(\frac{1}{1+\frac{\tau}{\epsilon}\left((H_{0}\frac{2\pi}{L})^{2}+k^{2}\right)}\left|\mathbf{q}\frac{2\pi}{L}\alpha^{(n-1)}_{\mathbf{q}}\right|^{2}+\frac{\left|\beta^{(n-1)}_{\mathbf{q}}\right|^{2}}{|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}+\frac{\epsilon}{\tau}}\biggr).$
Following the same argument as in (73), this simplifies to
@@
For the low-frequency terms, the independence of the particles $\widetilde{\mathbf{X}}_{1}^{(n)},\ldots,\widetilde{\mathbf{X}}_{P}^{(n)}\sim\mathbf{X}^{(n)}$ implies that
@@
Applying the attenuation factor $\frac{1}{|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}}$ and taking the expectation over the sum yields
@@
The factor $6H_{0}$ arises from a spherical shell summation: for each $1\leq h\leq H_{0}$, there are $6h^{2}$ frequency indices $\mathbf{q}$ with $L^{1}$-norm $|\mathbf{q}|_{1}=h$. The $h^{-2}$ scaling of the spectral terms $\frac{|\mathbf{q}\frac{2\pi}{L}|^{2}}{(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2})^{2}}$ cancels this growth, leaving a linear accumulation in $H_{0}$.
Combining the high-frequency error bound (84) and the low-frequency variance estimate (86) via the Parseval decomposition (79), we establish the desired final error bound.
The temporal error of the Euler-Maruyama semi-discretization over the time interval is stated as follows (see [milstein2004stochastic, p. 15-16]).
@@
Let $\{\mathbf{X}_{p}\}_{1\leq p\leq P}$ be $P$ independent realizations of the SDE (11), and let the Brownian increments $\mathbf{W}_{p}$ satisfy
@@
The Euler-Maruyama semi-discretization exhibits first-order strong convergence for additive noise, i.e., when the diffusion term is independent of the solution process. More precisely, for $t=n\tau\leq T$ with integer $n\geq 1$,
@@
where $C_{0}=C_{0}(J_{1},M_{1},M_{2},M_{3})$ is a positive constant independent of the time step $\tau$ and $t$.
The main convergence result is presented as follows.
@@
For the SIPF-PIC method with fourth-order particle-to-grid and second-order grid-to-particle interpolation, the mean absolute error at the final time T, defined as $\hat{d}^{(n)}:=\frac{1}{P}\sum_{1\leq p\leq P}|\mathbf{X}_{p}(n\tau)-\widehat{\mathbf{X}}_{p}^{(n)}|$, satisfies the following growth bound:
@@
$+\left.\left.C_{3}L^{2}H_{0}^{-2}+M_{3}L^{2}H^{-2}+2S_{1}M_{1}P^{-\frac{1}{2}}\right)\right),$
with probability at least $1-N_{T}e^{-\frac{S_{1}^{2}}{2}}$. Here, $\chi$ denotes the chemotactic sensitivity in the KS system (1), and $K$, $C_{0}$, $C_{1}$, $C_{2}$, $C_{3}$, and $S_{1}$ are positive constants derived in the proof, independent of the domain size $L$, grid resolution $H$, frequency cutoff threshold $H_{0}$, number of particles $P$, and time step $\tau$.
@@
First, we employ the discrete Gronwall inequality (Lemma 4.6) to derive an estimate for $d^{(n)}$. From Lemma 4.8, we have
@@
We decompose the gradient error term using the auxiliary-empirical and reference-auxiliary bounds from Lemmas 4.10 and 4.13, respectively. The gradient discrepancy is first split via the inequality as follows:
@@
Substituting the reference-auxiliary expectation bound (78) and the auxiliary-empirical bound (65) into the Gronwall framework yields the composite error growth constant $K=\sqrt{2TN_{2}}$ and residual term
@@
where the constants $C_{1}=513\pi^{3}\sqrt{N_{2}}$, $C_{2}=\sqrt{N_{2}}$, and $C_{3}=(2\pi)^{-2}\frac{\sqrt{N_{2}N_{1,2}}}{M_{0}}$ are defined in terms of the problem-dependent constants. The final error bound for $d^{\left(N_{T}\right)}$ follows from the estimate $(1+(K+\chi M_{2})\tau)^{N_{T}}\leq e^{T(K+\chi M_{2})}$ and the temporal accumulation of the residual terms, leading to the estimate as follows:
@@
which follows from the triangle inequality and the Cauchy-Schwarz inequality, combining with the temporal error bound from Lemma 4.15, we obtain the claimed convergence bound in (89).
@@
By optimizing the frequency cutoff threshold $H_{0}=H^{\frac{8}{13}}$, the error bound simplifies to
@@
with probability at least $1-N_{T}e^{-\frac{S_{1}^{2}}{2}}$.